Operation Summary for December 4, 2009 - December 7, 2009 |
The Integrated Luminosity for the period from 11/30/09 to 12/7/09 was 49.6 pb-1. Friday Dec. 4 The day shift began with Operations monitoring the Tevatron (TeV) store 7415 (40.87E30) and the Recycler stash (266.4E10), with the Antiproton Source stacking (24.5E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations terminated store 7415 at 1 PM and put the TeV into HEP shot setup. Operations established store 7416 at 3 PM with an average initial luminosity (AIL) of 249E30. Operators tuned the accelerators for increased intensities and efficiencies. Operations sent beam to the Meson MTest experiment when beam was available. During the evening shift, Operations disabled the MTest beam at 8 PM. Operations reported some downtime due to a Booster RF station (BRF6) tripping off on a bad flow indication; it reset. Saturday Dec. 5 The midnight shift began with Operations monitoring store 7416 (89.97E30) and the stash (230.9E10), with Pbar stacking (21.9E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7416 at 6 AM, and put the TeV into shot setup. Operations established store 7417 at 7:30 AM with an AIL of 255.4E30. During the day shift, Operations began sending beam to MTest at 8 AM. Operations reported at 1:19 PM that the 400 MeV area beam permit had tripped off due to a regulation problem with a gauss meter (L:SPEC) that started wandering. Linac personnel disabled the regulation loop for the power supply. During the evening shift, Operations disabled the MTest beam at 5:30 PM when a MTest power supply (F:MT5E) began tripping off on a ground fault indication; Operations called in EE Support personnel to check out the supply and replace the ground fault detector fuse (which tripped again by the end of the shift); troubleshooting continued. Operations terminated store 7417 at 8:29 PM and put the TeV into shot setup. Operations established store 7418 at 11:54 PM with an AIL of 272E30. Sunday Dec. 6 The midnight shift began with Operations monitoring store 7418 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, at 4:11 AM, Operations noted TeV proton beam problems that they correlated to the movement of a quadrupole magnet (C:D1Q3R). Operations reported that the MT5E problems continued – Switchyard personnel had operators set things up so the experiment could run without this power supply on. During the day shift, the MI low-level RF system required a reboot, which caused some downtime from 9:30 to 10 AM. Operations terminated store 7418 at 2:30 PM and then put the TeV into shot setup. Operations established store 7419 at 4 PM with an AIL of 277E30. Operations disabled the MTest beam at 2:30 PM; the experiment said they were finished for the day. During the evening shift, two Linac RF stations (LRF3 & 5) tripped off; they reset. Monday Dec. 7 The midnight shift began with Operations monitoring store 7419 (149.36E30) and the stash (152.1E10), with Pbar stacking (20.3E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Problems with MI beam positions required operators to drop the MI intensity while they tuned. Operations called in the Proton Source Specialist to assist Operations with troubleshooting and tuning beam coming out of the Linac and Booster. The MI Specialist also assisted in this tuning, but by phone. Operations terminated store 7419 at 6:04 AM and then put the TeV into shot setup. Operations established store 7420 at 7:36 AM with an AIL of 279.5E30. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow MI to conduct slipstacking studies. |
|
More Information Comments and Suggestions |
| last modified 12/07/2009 email Fermilab |
| Security, Privacy, Legal | |